# Thermal effects on warm chromoinflation

### v1

@@ III Thermal contributions to the background dynamics
$\bar{V}_{\rm eff}(\Psi)=\left(2+M_{\psi}^{2}\right)\frac{\Psi^{2}}{2}-\frac{2\xi}{3}\Psi^{3}+\frac{\Psi^{4}}{2},$ (40)
@@
$\Psi_{\pm}=\frac{1}{2}\left[\xi\pm\sqrt{\xi^{2}-2\left(2+M_{\psi}^{2}\right)}\right].$ (41)
@@
$\xi^{2}=\frac{9}{4}\left(2+M_{\psi}^{2}\right).$ (42)
This behavior of the effective potential, as seen here as function of the values of the parameters, resembles exactly that one for a first-order phase transition as a function of the temperature. The value of the temperature for which $\bar{V}_{\rm eff}(0)=\bar{V}_{\rm eff}(\Psi_{+})$ is satisfied would correspond to the critical temperature of the phase transition, as conventionally considered in the literature of Ginzburg-Landau phase transition when it is first order Goldenfeld:1992qy . We can here estimate this critical temperature if we use the slow-roll approximation obtained from the entropy evolution equation, Eq. (35),
$Ts\simeq\frac{\Upsilon}{3H}\dot{\phi}^{2}=\frac{\kappa T^{3}}{3fH}\dot{\phi}^{2},$ (43)
where we have used the expression (5) for the dissipation coefficient. Finally, from Eq. (LABEL:entropy), considering $g\psi\ll T$, and from the definitions given in Eq. (39), we obtain the result for the critical temperature for phase transition in terms of the parameters of the model,
@@
In warm inflation we have in general that $T/H$ increases during inflation Motaharfar:2021egj ; Das:2022ubr . Thus, after the critical temperature Eq. (LABEL:Tc) is reached, $\Psi_{+}$ becomes a local minimum. Hence, above $T_{c}$ it becomes energetically favored for the system to be driven to $\psi=0$. This happens rather fast, in less than one e-fold, as indicated by our numerical results, after which the gauge vacuum expectation value no longer plays a role in the dynamics. The above qualitative analytical analysis is confirmed by our numerical results that are presented in the next section.
IV Numerical Results
@@
Figure 3: The same as in Fig. 2, but for an initial value of the dissipation ratio is $Q_{0}=1.3\times 10^{-2}$.
@@
$V(\phi)=\Lambda^{4}\left[1+\cos(\phi/f_{D})\right],$ (45)
@@
where $\Lambda$ is the normalization of the potential (typically $\Lambda^{4}=m_{\phi}^{2}f_{D}^{2}$, where $m_{\phi}$ is the mass of the axion-like particle and $f_{D}$ its decay constant). Note that in standard cold natural inflation one typically requires $f_{D}\gtrsim M_{\rm Pl}/\sqrt{2}$ to have a slowly rolling regime for the inflaton Freese:2004un . To be concrete, we assume $\Lambda$ to correspond to some confinement energy scale of some UV complete theory. Inflation here would then happen at energy scales below $\Lambda$ but still larger than some IR energy scale for the confinement of the $SU(2)$ gauge field considered here, such that the presence of a thermal bath and the use of the dissipation coefficient Eq. (5) are justified.
@@
In warm inflation with a dissipation term like Eq. (5) it was also been shown Montefalcone:2022jfw that consistency of the natural inflation potential with the observations still requires $f_{D}\sim M_{\rm Pl}$. In particular, a more recent analysis Zell:2024vfn has also shown that this model can only sustain inflation consistent with the observations in the weak regime of warm inflation, in which case $Q=\Upsilon/(3H)\ll 1$. Motivated by these previous references, here we will not assume that $f$, which appears in the dissipation and chromoinflation expressions and that comes from the Chern-Simons interaction, and $f_{D}$, which appears in the potential are the same [^4] This can also be interpreted in terms of an effective value for the coupling $\lambda$, such that it would be analogous to consider a modified $\bar{\lambda}$ given by $\bar{\lambda}=\lambda f_{D}/f$.. This will give us some more freedom when playing with the parameters of the model [^5] It can be assumed in particular that $f_{D}$ is a much larger scale than $f$, which is motivated from recent works based on clockwork models Kaplan:2015fuy ; Choi:2015fiu .. By fixing $f_{D}$, $\lambda$ and $g$, we can obtain for instance $f$ by demanding to inflation to last a specific number of efolds for a given dissipation ratio $Q$.
@@
Figure 4: Snapshots of the effective potential Eq. (38) as a function of $\psi$ at specific values of e-folds. Panel (a) is in the absence of the thermal effects for $\psi$, while panel (b) shows the effects of including them.
@@
Figure 5: The global minimum of the effective potential Eq. (38) as a function of the temperature, for the cases of $Q_{0}=1.2\times 10^{-3}$ (panel a) and for $Q_{0}=1.3\times 10^{-2}$ (panel b).
@@
Specific examples are given in Figs. 2 and 3, where we show some of the relevant background quantities. In both examples shown in those figures the total number of e-folds of inflation was taken to be $N_{\rm infl}=55$ and for illustration purposes the parameters taken were $g=0.6$, $\lambda=100$, $f_{D}/M_{\rm Pl}=3/\sqrt{2}$ (as motivated e.g. from Ref. Montefalcone:2022jfw ). The initial dissipation ratio in Fig. 2 is $Q_{0}=1.2\times 10^{-3}$, from which the values of $f$ and $\Lambda$ are found to be $f\simeq 0.05M_{\rm Pl}$ and $\Lambda\simeq 0.003M_{\rm Pl}$, respectively. In Fig. 3 we have $Q_{0}\simeq 1.3\times 10^{-2}$, with $f\simeq 0.035M_{\rm Pl}$ and $\Lambda\simeq 0.0016M_{\rm Pl}$.
@@
The results displayed in Figs. 2 and 3 show that the most important effect of including the thermal corrections to the gauge field background is on the evolution of $\psi$ itself. In the absence of a thermal mass correction, the gauge field background is sustained throughout the inflationary evolution. However, in the presence of the thermal correction it is eventually driven to zero well before inflation ends. Furthermore, the larger is the initial value for the dissipation ratio $Q$, the sooner $\psi$ is driven to zero. For the parameters considered, this suppression of $\psi$ already happens in the weak regime of warm inflation, $Q\ll 1$. For the parameters considered and for a dissipation ratio $Q\gtrsim 5\times 10^{2}$, we find that the gauge field background already vanishes at the onset and remains null throughout the inflationary evolution.
@@
In principle, we could believe that the results could be changed by making $g$ very small, thus suppressing the thermal effects and allowing a nonvanishing value for $\psi$ to be sustained throughout inflation. However, this also affects the dissipation coefficient through its dependence on the coupling $g$, forcing either $\lambda$ to be larger or $f$ to be smaller. In general we find that larger values of $Q$ becomes hard to be obtained (we also recall the results of Refs. Montefalcone:2022jfw ; Zell:2024vfn of the difficulties of having warm inflation in the strong regime, $Q>1$, in axion-like inflaton potentials). In special we find that the larger is the dissipation ratio $Q$, one requires smaller couplings to support a gauge background $\psi\neq 0$.
@@
Another approach to favor chromoinflation in the current scenario would be to increase the gauge background field. This would require a gauge field mass, $m_{A}$, much larger than the temperature. This would suppress thermal effects from Eq. (27), as they would be Boltzmann-suppressed. However, this introduces a new challenge. A larger gauge field mass, resulting from a non-zero $\psi$, would also strongly suppress the dissipation coefficient in Eq. (5). This is analogous to what is expected to happen in the electroweak phase transition Cohen:1993nk : above the electroweak scale, massless gauge bosons allow rapid sphaleron processes; below the scale, mass acquisition suppresses these processes. Similarly, here, a larger gauge field mass would suppress the dissipation necessary for warm inflation. While cold chromoinflation could still occur, a warm inflation regime would be precluded.
@@
As discussed at the end of the last section, we can interpret the vanishing of $\psi$ as a true phase transition. This is explicitly illustrate in Fig. 4, where we show snapshots of the effective potential Eq. (38) as a function of $\psi$ at different values of e-folds for the case of $Q_{0}\simeq 1.3\times 10^{-2}$, which is the case shown in Fig. 3. From the results of Figs. 2 and 3 (e.g. from the panels (b) in those figures), this behavior of the minimum of the effective potential can be interpreted in terms of an increasing value of the temperature at the corresponding values of e-folds.
@@
The corresponding critical temperature for the (first-order) phase transition seen in Fig. 4(b) is $T_{c}\simeq 1.6\times 10^{-4}M_{\rm Pl}$, while for the case of the parameters shown in Fig. 1(d), i.e. for $Q_{0}=1.2\times 10^{-3}$ corresponds to $T_{c}\simeq 4.9\times 10^{-4}M_{\rm Pl}$. These values agree well with the simpler estimate given by Eq. (LABEL:Tc). Hence, the smaller is $Q_{0}$, the larger becomes $T_{c}$.
@@
The minimum of the effective potential, corresponding to the solution $\Psi_{+}$ in Eq. (41), as a function of the temperature, is displayed in Fig. 5. It shows the behavior typical of a order parameter as a function of the temperature when the phase transition is first order. The order parameter (which is here represented by $\psi_{+}$) jumps discontinuously from a nonnull to null value across the phase transition point.
V Conclusion
@@
In this paper, we investigated the impact of this thermal mass correction on the background gauge field $\psi$. Our findings reveal that the evolution of $\psi$ can be significantly influenced by thermal effects, particularly at larger dissipation ratios $Q$ characteristic of warm inflation. A non-vanishing initial value of $\psi$ can be rapidly driven to zero by these thermal effects. Subsequently, the dynamics effectively reduces to that of minimal warm inflation without a background gauge field. Conversely, a non-zero gauge field mass, induced by a non-vanishing $\psi$, suppresses sphaleron processes responsible for the dissipation term in Eq. (5). This suppression hinders warm inflation. Therefore, a successful warm inflation scenario in chromoinflation would be prevented.
@@
In this work, we have primarily focused on the background dynamics. It is crucial to extend this analysis to perturbations to understand how they differ from standard warm inflation scenarios without a background gauge field. The interplay between the background gauge field and thermal effects could lead to significant deviations. Furthermore, the backreaction effects, commonly studied in cold chromoinflation Maleknejad:2016qjz ; Maleknejad:2018nxz ; Ishiwata:2021yne , should be considered. These additional factors are essential for comparing the model’s predictions with observational data. We plan to address these aspects in a future work.
Acknowledgements.

### v2

@@ 3 Thermal contributions to the background dynamics
$\bar{V}_{\rm eff}(\Psi)=\left(2+M_{\psi}^{2}\right)\frac{\Psi^{2}}{2}-\frac{2\xi}{3}\Psi^{3}+\frac{\Psi^{4}}{2},$ (3.24)
@@
$\Psi_{\pm}=\frac{1}{2}\left[\xi\pm\sqrt{\xi^{2}-2\left(2+M_{\psi}^{2}\right)}\right].$ (3.25)
@@
$\xi^{2}=\frac{9}{4}\left(2+M_{\psi}^{2}\right).$ (3.26)
This behavior of the effective potential, as seen here as function of the values of the parameters, resembles exactly that one for a first-order phase transition as a function of the temperature. The value of the temperature for which $\bar{V}_{\rm eff}(0)=\bar{V}_{\rm eff}(\Psi_{+})$ is satisfied would correspond to the critical temperature of the phase transition, as conventionally considered in the literature of Ginzburg-Landau phase transition when it is first order [37]. We can here estimate this critical temperature if we use the slow-roll approximation obtained from the entropy evolution equation, eq. (3.19),
$Ts\simeq\frac{\Upsilon}{3H}\dot{\phi}^{2}=\frac{\kappa T^{3}}{3fH}\dot{\phi}^{2},$ (3.27)
where we have used the expression (2.5) for the dissipation coefficient. Finally, from eq. (LABEL:entropy), considering $g\psi\ll T$, and from the definitions given in eq. (3.23), we obtain the result for the critical temperature for phase transition in terms of the parameters of the model,
@@
In warm inflation we have in general that $T/H$ increases during inflation [18, 38]. Thus, after the critical temperature eq. (LABEL:Tc) is reached, $\Psi_{+}$ becomes a local minimum. Hence, above $T_{c}$ it becomes energetically favored for the system to be driven to $\psi=0$. This happens rather fast, in less than one e-fold, as indicated by our numerical results, after which the gauge vacuum expectation value no longer plays a role in the dynamics. The above qualitative analytical analysis is confirmed by our numerical results that are presented in the next section.
4 Numerical Results
@@
Figure 3: The same as in figure 2, but for an initial value of the dissipation ratio is $Q_{0}=1.3\times 10^{-2}$.
@@
$V(\phi)=\Lambda^{4}\left[1+\cos(\phi/f_{D})\right],$ (4.1)
@@
where $\Lambda$ is the normalization of the potential (typically $\Lambda^{4}=m_{\phi}^{2}f_{D}^{2}$, where $m_{\phi}$ is the mass of the axion-like particle and $f_{D}$ its decay constant). Note that in standard cold natural inflation one typically requires $f_{D}\gtrsim M_{\rm Pl}/\sqrt{2}$ to have a slowly rolling regime for the inflaton [39]. To be concrete, we assume $\Lambda$ to correspond to some confinement energy scale of some UV complete theory. Inflation here would then happen at energy scales below $\Lambda$ but still larger than some IR energy scale for the confinement of the $SU(2)$ gauge field considered here, such that the presence of a thermal bath and the use of the dissipation coefficient eq. (2.5) are justified.
@@
In warm inflation with a dissipation term like eq. (2.5) it was also been shown [40] that consistency of the natural inflation potential with the observations still requires $f_{D}\sim M_{\rm Pl}$. In particular, a more recent analysis [41] has also shown that this model can only sustain inflation consistent with the observations in the weak regime of warm inflation, in which case $Q=\Upsilon/(3H)\ll 1$. Motivated by these previous references, here we will not assume that $f$, which appears in the dissipation and chromoinflation expressions and that comes from the Chern-Simons interaction, and $f_{D}$, which appears in the potential are the same [^5] This can also be interpreted in terms of an effective value for the coupling $\lambda$, such that it would be analogous to consider a modified $\bar{\lambda}$ given by $\bar{\lambda}=\lambda f_{D}/f$.. This will give us some more freedom when playing with the parameters of the model [^6] It can be assumed in particular that $f_{D}$ is a much larger scale than $f$, which is motivated from recent works based on clockwork models [42, 43].. By fixing $f_{D}$, $\lambda$ and $g$, we can obtain for instance $f$ by demanding to inflation to last a specific number of efolds for a given dissipation ratio $Q$.
@@
Figure 4: Snapshots of the effective potential eq. (3.22) as a function of $\psi$ at specific values of e-folds. Panel (a) is in the absence of the thermal effects for $\psi$, while panel (b) shows the effects of including them.
@@
Figure 5: The global minimum of the effective potential eq. (3.22) as a function of the temperature, for the cases of $Q_{0}=1.2\times 10^{-3}$ (panel a) and for $Q_{0}=1.3\times 10^{-2}$ (panel b).
@@
Specific examples are given in figures 2 and 3, where we show some of the relevant background quantities. In both examples shown in those figures the total number of e-folds of inflation was taken to be $N_{\rm infl}=55$ and for illustration purposes the parameters taken were $g=0.6$, $\lambda=100$, $f_{D}/M_{\rm Pl}=3/\sqrt{2}$ (as motivated e.g. from ref. [40]). The initial dissipation ratio in figure 2 is $Q_{0}=1.2\times 10^{-3}$, from which the values of $f$ and $\Lambda$ are found to be $f\simeq 0.05M_{\rm Pl}$ and $\Lambda\simeq 0.003M_{\rm Pl}$, respectively. In figure 3 we have $Q_{0}\simeq 1.3\times 10^{-2}$, with $f\simeq 0.035M_{\rm Pl}$ and $\Lambda\simeq 0.0016M_{\rm Pl}$.
@@
The results displayed in figures 2 and 3 show that the most important effect of including the thermal corrections to the gauge field background is on the evolution of $\psi$ itself. In the absence of a thermal mass correction, the gauge field background is sustained throughout the inflationary evolution. However, in the presence of the thermal correction it is eventually driven to zero well before inflation ends. Furthermore, the larger is the initial value for the dissipation ratio $Q$, the sooner $\psi$ is driven to zero. For the parameters considered, this suppression of $\psi$ already happens in the weak regime of warm inflation, $Q\ll 1$. For the parameters considered and for a dissipation ratio $Q\gtrsim 5\times 10^{-2}$, we find that the gauge field background already vanishes at the onset and remains null throughout the inflationary evolution.
@@
In principle, we could believe that the results could be changed by making $g$ very small, thus suppressing the thermal effects and allowing a nonvanishing value for $\psi$ to be sustained throughout inflation. However, this also affects the dissipation coefficient through its dependence on the coupling $g$, forcing either $\lambda$ to be larger or $f$ to be smaller. In general we find that larger values of $Q$ becomes hard to be obtained (we also recall the results of refs. [40, 41] of the difficulties of having warm inflation in the strong regime, $Q>1$, in axion-like inflaton potentials). In special we find that the larger is the dissipation ratio $Q$, one requires smaller couplings to support a gauge background $\psi\neq 0$.
@@
Another approach to favor chromoinflation in the current scenario would be to increase the gauge background field. This would require a gauge field mass, $m_{A}$, much larger than the temperature. This would suppress thermal effects from eq. (3.11), as they would be Boltzmann-suppressed. However, this introduces a new challenge. A larger gauge field mass, resulting from a non-zero $\psi$, would also strongly suppress the dissipation coefficient in eq. (2.5). This is analogous to what is expected to happen in the electroweak phase transition [44]: above the electroweak scale, massless gauge bosons allow rapid sphaleron processes; below the scale, mass acquisition suppresses these processes. Similarly, here, a larger gauge field mass would suppress the dissipation necessary for warm inflation. While cold chromoinflation could still occur, a warm inflation regime would be precluded.
@@
As discussed at the end of the last section, we can interpret the vanishing of $\psi$ as a true phase transition. This is explicitly illustrated in figure 4, where we show snapshots of the effective potential eq. (3.22) as a function of $\psi$ at different values of e-folds for the case of $Q_{0}\simeq 1.3\times 10^{-2}$, which is the case shown in figure 3. From the results of figures 2 and 3 (e.g. from the panels (b) in those figures), this behavior of the minimum of the effective potential can be interpreted in terms of an increasing value of the temperature at the corresponding values of e-folds.
@@
The corresponding critical temperature for the (first-order) phase transition seen in figure 4(b) is $T_{c}\simeq 1.6\times 10^{-4}M_{\rm Pl}$, while for the case of the parameters shown in figure 2(d), i.e. for $Q_{0}=1.2\times 10^{-3}$ corresponds to $T_{c}\simeq 4.9\times 10^{-4}M_{\rm Pl}$. These values agree well with the simpler estimate given by eq. (LABEL:Tc). Hence, the smaller is $Q_{0}$, the larger becomes $T_{c}$.
@@
The minimum of the effective potential, corresponding to the solution $\Psi_{+}$ in eq. (3.25), as a function of the temperature, is displayed in figure 5. It shows the behavior typical of a order parameter as a function of the temperature when the phase transition is first order. The order parameter (which is here represented by $\psi_{+}$) jumps discontinuously from a nonnull to null value across the phase transition point.
5 Conclusion
@@
In this paper, we investigated the impact of this thermal mass correction on the background gauge field $\psi$. Our findings reveal that the evolution of $\psi$ can be significantly influenced by thermal effects, particularly at larger dissipation ratios $Q$ characteristic of warm inflation. A non-vanishing initial value of $\psi$ can be rapidly driven to zero by these thermal effects. Subsequently, the dynamics effectively reduces to that of minimal warm inflation without a background gauge field. Conversely, a non-zero gauge field mass, induced by a non-vanishing $\psi$, suppresses sphaleron processes responsible for the dissipation term in eq. (2.5). This suppression hinders warm inflation. Therefore, a successful warm inflation scenario in chromoinflation would be prevented.
@@
In this work, we have primarily focused on the background dynamics. It is crucial to extend this analysis to perturbations to understand how they differ from standard warm inflation scenarios without a background gauge field. The interplay between the background gauge field and thermal effects could lead to significant deviations. Furthermore, the backreaction effects, commonly studied in cold chromoinflation [20, 45, 46], should be considered. These additional factors are essential for comparing the model’s predictions with observational data. We plan to address these aspects in a future work.
Acknowledgments